Joe's Tire Company has two support departments, Personnel and Maintenance. The Maintenance Department costs of $80,000 are allocated on the basis of standard service hours used. The Personnel Department costs of $20,000 are allocated based on the number of employees. Costs of Departments A and B are $40,000 and $60,000, respectively.

Data on standard service hours and number of employees are as follows:

Maintenance
Dept.
Personnel
Dept.
Production Dept. A Production Dept. B
Standard service hours used 200 200 240 160
Number of employees 10 20 40 120


How much of the cost of the Personnel Department is allocated to Department B using the direct method?

a) $12,632

b) $15,000

c) $48,000

d) $8,000

e) $13,333
Answer #1

Personnel Department cost is $20,000

Allocation is based on the number of employees in production department

No of Employees In Production Dep A 40

No of Employees In Production Dep B    120

Total No of Employees In both Production Dep is = 40+120 = 160

Cost Allocation Between Production Dep A & B of Personnel Department $20,000

Production Dep A = (Cost X No of employees In Dep A)/Total No of Employees

Production Dep A = ($20,000X40)/160 = $5,000

Production Dep B= (Cost X No of employees In Dep B)/Total No of Employees

Production Dep B =($20000X120)/160= $15,000

Therefore From the options OPTION B is Correct